OpenStudy (lillian_a):

how many points of inflection are there for the function y=x+cos(2x) on the interval [0,pi]

OpenStudy (solomonzelman):

Note: There is a test that determines whether or not your inflection point is actually an inflection or not, but let's leave that for later.

OpenStudy (solomonzelman):

Do you have any problems with finding the second derivative?

OpenStudy (solomonzelman):

(( just making sure, is it calculus, or is it not? ))

OpenStudy (lillian_a):

yes it's calculus

OpenStudy (lillian_a):

its a calculator problem

OpenStudy (solomonzelman):

ti-84 calculator problem?

OpenStudy (lillian_a):

yes

OpenStudy (lillian_a):

second derivative: -4cos(2x)

OpenStudy (solomonzelman):

Oh, that was very good!

OpenStudy (solomonzelman):

\(\color{#000000 }{ \displaystyle f''(x)=-4\cos(2x) }\)

OpenStudy (solomonzelman):

We have a constraint: \(\color{#000000 }{ \displaystyle x \in [0,\pi] }\) (i.e. \(\color{#000000 }{ \displaystyle 0 \le x \le\pi }\)) So, set, \(\color{#000000 }{ \displaystyle 0=-4\cos(2x) }\) And tell me which x-solutions (if any) are within this interval

OpenStudy (solomonzelman):

(I don't recall any ti-calculator-functions that relate to finding infleection points, although we can do it with "normal" calculus)

OpenStudy (lillian_a):

how do i figure out the root between 0 and pi

OpenStudy (solomonzelman):

for what values of x is -4cos(2x)=0 ?

OpenStudy (solomonzelman):

can you list a couple of first positive values for which cos(x)=0?

OpenStudy (lillian_a):

Im not sure